# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v3

@@ 5.2 Performance evaluation
The reward in the RL algorithm is defined based on trip completion with additional damping redundancy terms. However, two scenarios with the same trip completion at the end of the simulation may exhibit distinct TTS, and the one with the lower value should be regarded to have demonstrated a superior overall performance. In addition, to better showcase antifragility and the associated performance loss, as represented by the shaded area in Fig. 13, the main performance indicator being evaluated in this work is the Total Time Spent (TTS), which is calculated by adding up the number of vehicles within the network at each second of the simulation.
Since urban road networks are always subject to capacity constraints, and as Sun et al. (2024) has recently proven the fragile nature of urban road networks in terms of progressive fragility, fully antifragile traffic control strategies may be impossible to achieve and will inevitably fall into being fragile when the magnitude of disruptions is large enough. Therefore, this paper aims to demonstrate that the proposed perimeter control algorithm is less fragile than the state-of-the-art baseline algorithms, and we normalize all the other perimeter control algorithms over the RL baseline method to study the relative antifragility. To quantify the progressive antifragility of different algorithms, we calculate the skewness of distribution based on the samples from the last 25 incremental episodes, with $\mu$ and $\sigma$ denoting the mean and the standard deviation:
@@
A skewness computed to be 00 means that the system itself or the applied algorithm makes the system neither fragile nor antifragile. On the other side, negative skewness indicates the distribution has a longer or fatter right tail and thus a higher degree of concavity in the performance function, which showcases antifragility, whereas a positive skewness signifies fragility. Therefore, other than demonstrating a superior performance regarding a lower TTS, the proposed method should also showcase a lower skewness than the baseline method.
@@ 6 Results
As mentioned in the previous section, there are 8 simulation scenarios based on the combination of demand/supply disruption, static/incremental disruption, and with/without uncertainty, with each one containing the performance of 4 perimeter control algorithms to be tested: NC, MPC, the baseline RL algorithm (Zhou and Gayah, 2021), and our proposed RL-based algorithm. Here, we first present the results of simulations with a static magnitude of disruptions to showcase proto-antifragility, followed by the results with incremental magnitudes to demonstrate progressive antifragility. Each scenario is tested for 25 simulation iterations.
6.1 Proto-antifragility
Systems with the property of proto-antifragility can learn from past adverse events and anticipate possible ongoing disruptions of similar magnitudes to enhance future performance. To validate such property for the RL-based algorithms, after training the agent with the base demand profile for 50 episodes, we apply disruptions with static magnitude for another 25 episodes.
Fig. 7 demonstrates the performance curves of the four algorithms under no uncertainty or with uncertainty. Each data point on any performance curve for a particular episode represents the TTS averaged from 25 simulation iterations of 3 hours. The shaded area represents the performance variance due to the innate randomness in the training process of the RL agents or combined with uncertainties introduced following the shuffling scheme described in Section. 5.1.
(a) Without disruption uncertainty
(b) With disruption uncertainty
Figure 7: Performance curves under static demand disruptions
In the 50 episodes of training with the base demand profile in Fig. 7(a), the proposed RL algorithm outperforms both NC and the benchmark RL algorithm and demonstrates performance comparable to that of MPC. In the following 25 episodes with disruption, both RL-based algorithms exhibited improved performance following an initial rise of TTS at the transition and outperform NC and MPC. Overall, the proposed RL algorithm performs better than the baseline RL algorithm before the introduction of the static disruption and shows an even superior performance of about $4.2\%$ after the disruption comes into play. In addition, the performance variance of the proposed RL algorithm is significantly smaller than the benchmark RL algorithm. In Fig. 7(b), as disruption magnitude is set to be stochastic following a Gaussian distribution, the performance of NC and MPC also contains a shaded area. The shaded area is rather stable under uncertainty due to the reshuffling mechanism explained in Section 5 to reshuffle the uncertainty multiplier list $[\epsilon_{i}]$ by $1$ after each iteration. Despite the enlarged variance for both RL-based methods, the averaged TTS curves are largely the same as in 7(a) with no uncertainty. The performance gain of the proposed RL algorithm compared to the RL baseline under disruption uncertainty is $3.9\%$, marginally smaller compared to the performance without uncertainty.
Likewise, as demonstrated in Fig. 8, similar results can be observed in the study of static supply disruption. The averaged TTS curves show a $3.0\%$ performance gain of the proposed RL algorithm compared to the baseline RL algorithm under no disruption uncertainty. In Fig. 8(b), the performance gain is marginally higher and comes to $3.5\%$ under the presence of disruption uncertainty. However, the performance variance of the proposed algorithm in Fig. 8(a) is greater with disruption uncertainty compared to the baseline RL algorithm, but still, its shaded area in blue lies predominantly under the gray shaded area, indicating the superior performance of the antifragile RL algorithm.
(a) Without disruption uncertainty
(b) With disruption uncertainty
Figure 8: Performance curves under static supply disruptions
6.2 Progressive antifragility
Progressive antifragility describes a system that exhibits a nonlinear response to growing disruptions. Since an increase in traffic density or vehicle accumulation within a network will inevitably reduce average vehicle speed or trip completion and increase TTS, a concave response between TTS and vehicle accumulation can indicate a progressively antifragile system. Instead of 25 episodes of a static disruption, a linearly increasing magnitude of demand or supply disruption is simulated, and the distribution skewness is calculated to determine system (anti-)fragility.
@@ 6.2.1 Demand disruption
In Fig. 9(a), when no disruption uncertainty is present in the last 25 episodes of incremental disruptions, the proposed RL algorithm performs the best out of all four tested algorithms. The shaded area represents the variance from the 25 iterations of the simulation and is lower than the RL baseline. Furthermore, the performance curve of the proposed method also seems less convex than the other algorithms, demonstrating its relative antifragility. Before demonstrating antifragility quantitatively, it should be noted that the TTS curves from NC and MPC are only partially illustrated, as their simulated network came into gridlock quicker when faced with a major disruption with the vehicle accumulation reaching maximal accumulation. And when the network is fully gridlocked, it cannot recover to an uncongested state even when the surging demand has declined. In this case, TTS will only grow linearly with the magnitude of distribution and lose the ability to showcase skewness, which will be further detailed below in Fig. 9(c).
@@
(b) Normalized over Zhou & Gayah (2021)
@@
(b) Normalized over Zhou & Gayah (2021)
@@
In Fig. 9(b), the performance difference normalized over the baseline RL algorithm is illustrated with a polar plot. Only the episodes $50-75$ under incremental demand are shown. Instead of TTS variance, the shaded blue area highlights the performance gain of our proposed method compared to the state-of-the-art baseline RL algorithm. As observed, the TTS curves for NC, MPC, and the proposed RL algorithm all diverge from the baseline method at an accelerating rate, indicating diverse degrees of relative (anti-)fragility. The main comparison here is the two RL-based algorithms, and our proposed algorithm shows an ultimate $9.3\%$ performance gain after 25 episodes with incremental demand disruption. Also, the disruption magnitude between the 62-rd and 63-rd episodes under incremental disruption is about the same as in the simulation with static disruption, and it is interesting to see that the performance gain here is approximately $5.2\%$, significantly higher than the $4.2\%$ improvement with a static magnitude of disruption, hinting the proposed method adapts relatively faster under an incremental environment.
To analyze the antifragile properties of each perimeter control algorithm, we compute the distribution skewness for each method up to the $n$-th episode, as in Fig. 9(c), so that the skewness value at the final episode reflects the distribution skewness across all $25$ episodes with incremental disruptions. Since calculating skewness requires a minimum sample size to yield meaningful results, a buffer of $5$ episodes is included at the start of the incremental disruption phase. Initially, the skewness for all four methods starts at a low value. However, the skewness for NC and MPC increases rapidly to approximately $1$, whereas the skewness for the two RL-based algorithms rises more slowly. Among these, the skewness curve in blue of the proposed algorithm lies mostly below the baseline curve in gray, particularly when the magnitude of disruption is high, achieving the lowest final skewness at $0.29$. It should be noted that the results from both NC and MPC indicate that these two algorithms render the network quickly into gridlock under the presence of demand disruptions, where the TTS will only grow linearly with the magnitude of disruptions and yield a skewness of 00. Therefore, we use dashed lines to represent the maximal skewness if the skewness of a particular algorithm has ever plateaued.
When uncertainty is introduced to the magnitude of disruptions, as shown in Fig. 10, the results remain largely consistent with the non-stochastic scenario, with the performance curves being less smooth, suggesting that the presence of uncertainty adds complexity to the training process. Compared to the baseline algorithm, the proposed RL algorithm demonstrates a continuously increasing performance gain, as shown in Fig. 10(b), reaching $9.2\%$ when the magnitude of disruption becomes significant. The unevenness in performance is also reflected in the value of skewness in Fig. 10(c), despite the final skewness of the two RL-based algorithms $0.32$ and $0.29$ being almost the same $0.32$ and $0.28$ as in Fig. 9(c). The final skewness values for NC and MPC are noticeably lower. The introduction of uncertainty causes gridlock to occur sooner, resulting in the premature entering into plateauing for both the NC and MPC, particularly when considering the convex fragile performance of the network.
@@ 6.2.2 Supply disruption
TTS curves of the four studied algorithms under deterministic supply disruption are shown in Fig. 12. The magnitude of incremental supply disruption is tuned so that the performance curve from NC is largely the same as with incremental demand disruption. However, the TTS from MPC and RL-based algorithms has somewhat increased, showing an increase of roughly $15\%$. The difference in performance has demonstrated the distinct levels of difficulty demand and supply disruption have posed for the learning of the RL agents. The relatively faster-growing skewness in Fig. 11(c) is yet another evidence of such challenges. The state of the baseline RL algorithm is composed of $[n_{ij}(t),q_{ij}(t)]$, whereas it is $[n_{ij}(t),\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ in the proposed algorithm. Although the baseline method relies on additional information of $[q_{ij}(t)]$, it becomes misinformation under demand disruption, whereas the derivatives $[\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ have a similar function as $[q_{ij}(t)]$ but being less susceptible to demand disruption. When it comes to disruption on the supply side, the situation differs because neither of the state definitions of the two RL-based algorithms is capable of accurately reflecting the model shifting. Still, with a faster changing rate of the derivatives, the agents in the proposed algorithm can make better decisions and gradually adapt to the new model. As shown in Fig. 11(b), the performance gain can grow up to $12.2\%$ when the supply disruption is substantial with $\lambda$ being $0.145$, demonstrating the superior performance of the proposed method compared to the baseline method. Furthermore, in Fig. 11(c), the blue skewness curve lies constantly below the baseline curve in gray, with an ultimate skewness of $0.41$, which indicates that the proposed RL algorithm is relatively more antifragile than the baseline algorithm.
@@
(b) Normalized over the baseline RL algorithm
@@
(b) Normalized over the baseline RL algorithm
@@
When supply disruption uncertainty is introduced into the simulation, as shown in Fig. 12, the results are similar to those observed under demand disruption with uncertainty. First, the TTS difference is $11.8\%$, showing minimal variation between scenarios with and without disruption. The skewness of the two RL-based algorithms remains almost unchanged, shifting only slightly from $0.41$ and $0.52$ to $0.42$ and $0.51$ for the proposed and baseline algorithms, respectively. In contrast, the skewness values for NC and MPC decrease from $0.88$ and $0.78$ to $0.82$ and $0.73$, reflecting the fact that the traffic state has already reached the gridlock accumulation without uncertainty, and by the incorporation of uncertainty, the network is prone to get gridlocked even more easily. Likewise, even with an average of 25 simulation iterations, the performance curves under supply disruption with uncertainty are less smooth, which is a similar pattern under demand disruption uncertainty.

### v4

@@ 5.2 Performance evaluation
The reward in the RL algorithm is defined based on trip completion with additional damping and redundancy terms. However, two scenarios with the same trip completion at the end of the simulation may exhibit distinct Total Time Spent (TTS), and the one with the lower value should be regarded as having demonstrated a superior overall performance. In addition, to better showcase antifragility and the associated performance loss, as represented by the shaded area in Fig. 16, the main performance indicator being evaluated in this work is the TTS, which is computed by adding up the number of vehicles within the network at each second of the simulation.
Since urban road networks are always subject to capacity constraints, and as Sun et al. (2024) has recently proven the fragile nature of urban road networks, fully antifragile traffic control strategies may be impossible to achieve and will inevitably fall into being fragile when the magnitude of disruptions is large enough. Therefore, this paper aims to demonstrate that the proposed perimeter control algorithm is less fragile than the state-of-the-art baseline algorithms, and we normalize all the other perimeter control algorithms over the RL baseline method to study the relative antifragility. To quantify antifragility of different algorithms, we calculate the distribution skewness based on the samples from the last 25 incremental episodes, with $\mu$ and $\sigma$ denoting the mean and the standard deviation:
@@
A skewness computed to be 00 means that the system itself or the applied algorithm makes the system neither fragile nor antifragile. On the other side, negative skewness indicates the distribution has a longer or fatter right tail and thus a higher degree of concavity in the performance function, which showcases antifragility, whereas a positive skewness signifies fragility. For the detailed methodology regarding skewness and antifragility, interested readers may refer to Sun et al. (2024). Therefore, other than demonstrating a superior performance regarding a lower TTS, the proposed method should also showcase a lower skewness than the benchmarks.
@@ 6 Results
Following the simulation setup in Section 5, we first evaluate the performance and antifragile characteristics of the studied approaches under idealized conditions with full observability. Subsequently, we examine how the RL-based algorithms perform under real-world constraints with limited observability. Each scenario is evaluated over 25 simulation runs. Performance is also assessed with and without stochastic variations for both full and limited observability settings.
6.1 Performance under full observability
Under idealized full observability, all algorithms are given perfect information on $[n_{ij}(t),q_{ij}(t)]$. As discussed in Section 5, MPC-MHE is only applied under supply disruption, where a linearly increasing reduction factor $r$ is assumed and can be estimated through MHE.
@@ 6.1.1 Demand disruption
Demand disruption can be understood as a surging traffic demand within a short period, as illustrated in Fig. 7(a) on top of $q_{22,N}$ as a critical scenario. With Fig. 8(a) we present the performance curves of different studied algorithms, i.e., MPC as in Geroliminis et al. (2013) in orange, baseline RL in Zhou and Gayah (2021) in gray, and the proposed antifragile RL algorithm in blue under the cordon-shaped perimeter control environment. The first 50 episodes represent the training process for the RL-based algorithms under a static demand profile, which follows a normal distribution. The TTS curves first drop down and reach a comparable performance as MPC, corresponding to the conclusion in Zhou and Gayah (2021). The 50 training episodes are followed by another 25 episodes with linearly increasing demand disruption, with no stochasticity assumed on top of the linearly growing magnitude. The studied algorithms showcase distinct capabilities of learning and adapting to such disruptions, and the performance variance from the 25 iterations of the simulation. It is obvious that the proposed antifragile RL algorithm exhibits both superior performance and reduced variance, as indicated by the significantly narrower shaded area of its curve. Furthermore, the performance curve of the proposed method also seems less convex than the other algorithms, demonstrating its relative antifragility. Note that due to the inherent variability of RL-based algorithms as well as the high sensitivity of the distribution skewness, which will be computed below, on such variability, we apply a sliding window of 5 episodes to smooth the performance curve.
Such superior performance is further quantitatively illustrated with a polar plot in Fig. 8(b), which shows the performance difference in TTS reduction normalized over the baseline RL algorithm. Only the episodes $50-75$ under incremental demand are shown. Instead of variance, the shaded blue area highlights the performance gain of our proposed method compared to the baseline RL algorithm, showing a steady increase in TTS reduction and being largely concave. The average performance gain is $9.3\%$reaching $21.0\%$ after 25 episodes under incremental demand disruption. On the other hand, although MPC performs relatively better when the magnitude of disruption is low, its performance quickly draws near to the baseline RL under high demand disruptions, achieving only a $4.8\%$ improvement by the end of the simulation.
To quantify the antifragile properties of each perimeter control algorithm, we compute the distribution skewness for each method up to the $n$-th episode, as in Fig. 8(c), meaning the value of the skewness curve at the final episode reflects the antifragility characteristics across all $25$ episodes under incremental disruptions. Since the calculation of skewness requires a minimum sample size to yield meaningful results, a buffer of $5$ episodes is included at the start of the incremental disruption phase, i.e., episode $50-55$. Initially, the skewness for all three methods starts at a low value. However, the skewness of the baseline RL in gray rises rapidly, whereas the skewness of the MPC algorithm progresses even faster to $0.86$, despite the good performance when the magnitude of disruption is low, as depicted in Fig. 8(c). The skewness of the antifragile RL-based algorithm in blue, on the contrary, increases more slowly, and lies largely below the baseline curve, especially when the magnitude of disruption is high, ultimately achieving the lowest final skewness at $0.46$.
@@
(b) Normalized over baseline RL
@@
(b) Normalized over baseline RL
@@
Fig. 9 shows the performance when stochasticity is introduced to the magnitude of disruptions, following a disruption shuffling scheme discussed in Section 5.1 to ensure each episode between $50-75$ goes through the same total magnitude of disruptions. The results display a consistent trend with the non-stochastic scenario, with the performance curves being less smooth, as in Fig. 9(b), suggesting that the presence of uncertainty adds complexity to the training process. The performance of both RL-based methods deteriorates slightly, but still, the performance of the proposed antifragile RL algorithm outperforms the baseline RL by an average of $8.0\%$ and achieves $18.6\%$ performance gain at the end of the simulation under incremental demand disruptions, as well as a gain of $12.9\%$ compared to MPC. The fact that the TTS difference curve is approximately convex for MPC and concave for the antifragile RL algorithm can be quantified in Fig. 9(c). While the final distribution skewness of the baseline and the antifragile RL algorithms is $0.70$ and $0.56$, respectively, the skewness of MPC records a value of $0.93$, indicating its high fragility under incremental demand disruptions. An interesting phenomenon worth noticing is the spike of the skewness curve of the proposed algorithm around episode $60-65$, which can be observed in not only Fig. 8(c) and 9(c), but also in the results from supply disruption in the following discussion. The observed effect is likely due to the interplay between multiple RL hyperparameters, including the memory buffer, step size, and learning rate. After episode $50$, the replay buffer contains a mixture of training data under both disrupted and undisrupted scenarios, resulting in learning from a mixed dataset and causing a moderate deviation from optimal performance.
@@ 6.1.2 Supply disruption
Supply disruption can be modeled as a proportional reduction in both the capacity and the maximum vehicle accumulation on the MFD profile, as represented in Fig. 7(b). As discussed in Section 5.1, only the MFD profile of the inner region is reduced as a critical scenario in our simulation, where the outer region remains intact. Figure 11 presents the performance curves of the studied algorithms under supply disruption. In addition to MPC (orange), baseline RL (gray), and the proposed antifragile RL (blue), the MPC-MHE variant (brown) is also included. By modeling the supply disruption magnitude coefficient $r$ as an unknown system parameter, the MHE module estimates $r$ in real time to capture the extent of the MFD reduction. Fig. 10(a) shows the performance curves of the four studied algorithms under the no supply disruption for the first $50$ episodes, followed by linearly declining MFD during episode $50-75$. Unlike the performance under demand disruption in Fig. 8(a), where MPC performs better at lower disruption magnitudes, the TTS curve of the proposed antifragile RL algorithm remains consistently similar or below those of the other methods, indicating superior performance across the entire range of disruption.
@@
(b) Normalized over baseline RL
@@
(b) Normalized over baseline RL
@@
The performance gain is quantitatively analyzed in Fig. 10(b), which shows an average TTS reduction of $16.7\%$ from the proposed antifragile RL compared to the baseline RL algorithm under incremental supply disruption, and reaches an ultimate performance improvement of $39.4\%$. It is worth noting that the MPC-MHE curve in brown also delivers strong performance, achieving a $31.4\%$ TTS reduction at episode $75$, highlighting the effectiveness of incorporating the supply disruption magnitude coefficient $r$ to be estimated within the MHE framework. Still, the performance does not quite match that of the antifragile RL algorithm under high magnitudes of supply disruption. Moreover, MHE relies on accurately modeling both the system dynamics and the evolution of the model itself, which is a highly challenging requirement in real-world scenarios. Another shortcoming of MHE is the marginally worse performance when no supply disruption is present, i.e., during episode $0-50$, showing a $4.0\%$ reduction compared to MPC.
The distribution skewness up to the $n$-th episode under disruption is demonstrated in Fig. 10(c). The proposed antifragile RL method exhibits noticeably lower skewness across episodes $50-75$ compared to the other approaches, with a final skewness of $0.49$, whereas the skewness of the other three methods hovers around $1$. As discussed in Section 6.1.1, the spike in the skewness curve of the antifragile RL algorithm around episode $60-65$ may be caused due to the mixture of training data under both disrupted and undisrupted scenarios. Note that the baseline RL algorithm exhibits a plateau in skewness toward the end of the simulation, indicating the network has already succumbed to a gridlock. When a network is gridlocked, TTS increases linearly with the current vehicle accumulation. This implies that beyond a certain threshold, demand disruption will only lead to a linear TTS growth instead of exponential, while supply disruptions primarily accelerate the onset of gridlock without significantly affecting the eventual TTS. As a result, when overreaching the disruption threshold, the skewness of TTS initially plateaus and then gradually approaches zero. Such a phenomenon is more pronounced in the below Fig. 11(c). Since the skewness value being zero suggests the system is neither fragile nor antifragile, the proposed algorithm is applicable under typical disruptions through antifragile control, while not designed to handle catastrophic scenarios where the network rapidly descends into gridlock.
When supply disruption uncertainty is introduced into the simulation, as shown in Fig. 11, the general trend is similar to that observed under demand disruption with uncertainty. The performance curves of MPC, MPC-MHE, and the antifragile RL algorithm all deteriorate in the presence of stochasticity, while the baseline RL algorithm surprisingly shows improved performance. As discussed earlier, this improvement can be attributed to the onset of gridlock and its effect on supply disruptions. When stochasticity amplifies the supply disruption magnitude coefficient, its impact is minimal since the network is already gridlocked. In contrast, when stochasticity attenuates the coefficient, it reduces TTS and enhances performance. As a result, the overall performance of the baseline RL algorithm appears to improve. Nevertheless, the performance gain of the proposed algorithm is demonstrated quantitatively in Fig. 11(b), showing an average improvement of $14.6\%$ and reaching $27.1\%$ at the final episode, while MPC and MPC-MHE show an ultimate $6.4\%$ and $19.1\%$ performance gain, respectively. 11(c) summarizes the skewness of the studied algorithms, with our proposed antifragile RL achieving the lowest value of $0.70$. Although the ultimate distribution skewness of the baseline RL reaches $0.79\%$, its peak skewness is $0.90\%$ at episode $71$.
6.2 Performance under real-world limited observability
As discussed in Section 4.1, obtaining real-world measurements of $n_{ij}$ and $q_{ij}$ is highly challenging. Therefore, based on Eq. 8 and 9, we evaluate the performance of the baseline and proposed antifragile RL-based algorithms under conditions of real-world limited observability. It is important to highlight that the performance of the proposed algorithm under limited observability is not directly comparable to MPC or MPC-MHE under full observability, given their fundamentally different observability requirements. Fig. 13 and 13 illustrate the performance of algorithms in the presence of linearly increasing demand disruptions, modeled either deterministically or with disruption magnitude uncertainties.
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 12: Performance curves under incremental demand disruptions
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 14: Performance curves under incremental supply disruptions
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
In Fig. 12(a), while the TTS curve of the baseline RL algorithm in light coral is approximately the same under both full and limited observability, the performance curve of the antifragile RL algorithm under limited observability in light teal falls between that of the baseline RL and the antifragile RL with full observability. This indicates that the algorithm sacrifices some performance in exchange for depending on more realistic observable measurements. The TTS difference in Fig. 12(b) marks the performance difference between the baseline and antifragile RL-based algorithms under limited observability, taking an average of $6.1\%$ performance improvement and reaching $15.0\%$ at the end of the simulation. It should be highlighted that the performance trade-off to achieve realistic observability, i.e., the performance difference between limited (light teal) and full observability (blue) of the proposed algorithm, is an average of $4.0\%$. The distribution skewness is quantitatively illustrated in Fig. 12(c). Although taking a higher skewness value at the onset of disruption, the proposed method yields a rather low ultimate skewness of $0.49$ under limited observability, lying between the skewness observed with full observability ($0.46$) and that of the baseline RL under limited observability ($0.62$).
When under demand disruption with magnitude stochasticity as shown in Fig. 13(a), the proposed RL algorithm suffers from slightly larger deterioration compared to the baseline RL, mainly because the performance of the baseline RL is close to network gridlock. Still, the TTS difference in Fig. 13(b) showcases the superiority of the proposed algorithm with an average of $4.2\%$ performance gain and reaches $8.2\%$ at the final episode. The performance loss for the trade-off of observability is $4.3\%$. The skewness under disruption stochasticity exhibits some volatility, which shows a value of $0.61$ and is identical to that of the baseline RL under limited observability.
In Fig. 15 and 15, the performance of both algorithms under supply disruptions with limited observability follows similar trends to those under demand disruption. The shaded area in Fig. 13(b) exhibits the TTS difference between the two algorithms with limited observability, which yields an average performance gain of $11.1\%$ and achieves a final value of $30.7\%$. The observability trade-off is $5.5\%$, taking an even smaller percentage of the average gain than under demand disruption. Likewise, its skewness $0.70$ at episode $25$ is also positioned between the baseline RL with limited observability and the antifragile RL with full observability. When under disruption magnitude stochasticity in Fig. 15(a), as explained, the performance of the baseline RL method improves at high disruption magnitude, because the network is close to gridlock. When stochasticity brings in higher and lower magnitudes of disruptions, stronger disruption has little effect on the performance, whereas weaker disruption leads to better performance. An intuitive proof can be observed at episode 70, where the TTS of the baseline RL both with and without uncertainty in Fig. 14(a) and Fig. 15(a) is around $1\cdot 10^{8}$ seconds, indicating the gridlock taking place shortly afterwards. Fig. 15(b) exhibits an ultimate performance gain of $12.5\%$. Although this scenario shows the highest observability trade-off of $10.1\%$, the skewness of the proposed antifragile RL is rather low with a value of $0.69$ compared to the other algorithms.